# Towards Optimal Environmental Policies: Policy Learning under Arbitrary Bipartite Network Interference

### v2

@@ Abstract
The substantial effect of air pollution on cardiovascular disease and mortality burdens is well-established. Emissions-reducing interventions on coal-fired power plants—a major source of hazardous air pollution—have proven to be an effective, but costly, strategy for reducing pollution-related health burdens. Targeting the power plants that achieve maximum health benefits while satisfying realistic cost constraints is challenging. The primary difficulty lies in quantifying the health benefits of intervening at particular plants. This is further complicated because interventions are applied on power plants, while health impacts occur in potentially distant communities, a setting known as bipartite network interference (BNI). In this paper, we introduce novel policy learning methods based on Q- and A-Learning to determine the optimal policy under arbitrary BNI. We derive asymptotic properties and demonstrate finite sample efficacy in simulations. We apply our novel methods to a comprehensive dataset of Medicare claims, power plant data, and pollution transport networks. Our goal is to determine the optimal strategy for installing power plant scrubbers to minimize ischemic heart disease (IHD) hospitalizations under various cost constraints. We find that annual IHD hospitalization rates could be reduced in a range from 20.66–44.51 per 10,000 person-years through optimal policies under different cost constraints.
Keywords: cost-constrained policy learning, environmental policy, power plants
@@ 1 Introduction
From a statistical perspective, policy learning methods for emissions-reducing policies must accommodate two key problem features that differ significantly from more conventional causal inference settings. First, intervention strategies are implemented in coal-fired power plants (“intervention units”), but health impacts are measured in surrounding communities (“outcome units”). This creates a so-called bipartite data structure. Second, the complex process in which airborne pollutants react in the atmosphere and are transported (e.g., by the wind) means that intervening at a single power plant can potentially affect health in many distant communities, resulting in a network of connections between power plants and proximal communities. In causal inference, the presence of such connections creates interference, where one unit’s treatment can spill to (possibly many) untreated units and can affect their potential outcomes. The aggregate of these two features has been described as bipartite network interference (BNI) (Zigler & Papadogeorgou, 2018). Although a small number of papers have proposed approaches for conducting causal inference under BNI, optimal policy learning methods under BNI have not been explored to date.
@@ 1.1 Related Work
The literature on causal analysis under BNI is relatively sparse. Here we highlight primary works in the field and related works in non-BNI settings. The causal BNI literature began with Zigler & Papadogeorgou (2018), introducing estimators for causal effects under partial interference, adapting standard inverse probability weighted estimators to this simplified setting. The BNI literature expanded to accommodate more general causal estimands under experimental (Pouget-Abadie et al., 2019; Doudchenko et al., 2020) and observational settings (Zigler & Papadogeorgou, 2018; Zigler et al., 2023; Chen, Bargagli-Stoffi, Kim & Nethery, 2024), including average and heterogeneous treatment effects. More recent work considers the estimation of causal effects from longitudinal data (Chen, Bargagli-Stoffi, Kim, Henneman & Nethery, 2024; Song & Papadogeorgou, 2024). To the best of our knowledge, no work has considered the problem of policy learning in the context of BNI. The closest related works consider policy learning under interference in non-BNI settings (Zhang & Imai, 2024; Su et al., 2019; Viviano, 2024; Viviano & Rudder, 2020). Other works have considered policy learning problems under safety or budget constraints, but operate under non-BNI settings (Qiu et al., 2022; Laber et al., 2018; Xu et al., 2024). It is worth noting that with some work, one can use plug in estimates from currently established work on causal estimation (Zigler & Papadogeorgou, 2018; Doudchenko et al., 2020) in order to estimate treatment effects and optimal policy strategies under BNI. However, those methods are either not developed for observational studies, unable to rigorously accommodate arbitrary interference, and/or lack double-robustness properties for targeted policy recommendations.
@@ 1.2 Contributions and Organization
In this paper, we develop and apply methods to learn optimal power plant scrubber installation policies to minimize hospitalization burdens for ischemic heart disease (IHD) under cost constraints. Our approach will leverage a rich dataset of Medicare claims, power plant locations and features, power plant pollution transport networks, community features, and scrubber installation costs for the eastern U.S. Building methodologically on the work of Robins (2004), we propose Q-Learning and A-Learning methods for policy learning under arbitrary BNI. In contrast to much of the existing work on interference, our approach enables estimation and inference for the average treatment effects of intervening on intervention units without the need for (1) clustering assumptions or (2) causal estimands that might suffer of subjective choices of the exposure mapping such as spillover or direct effects (Hudgens & Halloran, 2008; Bargagli-Stoffi et al., 2024). We derive the asymptotic properties of these methods and demonstrate their finite sample performance via simulation studies, finding robustness of A-Learning to model misspecification in terms of estimation and inference, compared to Q-Learning. We use the proposed methods to determine the power plant scrubber installation strategy that would optimally reduce Ischemic Heart Disease (IHD) hospitalization outcomes within the United States, under cost constraints, finding that annual IHD hospitalizations could be reduced in a range from 13,353 (20.66 per 10,000 person-years) to 31,906 (44.51 per 10,000 person-years) based on different financial budget constraints.
The rest of the paper is organized as follows: Section 2 introduces our Q- and A-Learning methods and their theoretical properties. Section 3 details an extensive semi-simulated Monte Carlo study carried out to assess performance of the methods in finite samples. In Section 4, we present the details and results of our scrubber installation policy analysis. Finally, we end with discussion and limitations in Section 5. Proofs are deferred to Supp. A.
@@ 2.1 Setup and Objective
Suppose we have $J$ intervention units indexed by $j\in[J]$ and $n$ outcome units indexed by $i\in[n]$. Let $\mathbf{X}^{out}_{i}$ denote the vector of covariates corresponding to outcome unit $i$, and $\mathbf{X}^{int}_{j}$ denote the vector of covariates corresponding to intervention unit $j$. $\mathbf{\mathbb{X}}_{1:J}$ will denote the covariates for all intervention units or $\mathbf{\mathbb{X}}_{1:J}=\{\mathbf{X}^{int}_{j}\}_{j\in[J]}$. $\mathbf{H}\in\mathbb{R}^{n\times J}$ or $\{\mathbf{H}_{ij}\}$ denotes the ‘interference map’ or bipartite adjacency matrix, which we will take to be known. The elements of $\mathbf{H}$ correspond to the strength of pollution transport from a particular intervention unit $j$ to each outcome unit $i$. $\mathbf{H}_{j}^{\top}=(\mathbf{H}_{1j}\dots\mathbf{H}_{ij}\dots\mathbf{H}_{nj})^{\top}$ denotes a particular column of the interference map. Similarly, $\mathbf{H}_{i}=(\mathbf{H}_{i1}\dots\mathbf{H}_{ij}\dots\mathbf{H}_{iJ})$ denotes a particular row of the interference map—that is, $\mathbf{H}_{i}$ represents the strength of connections between a particular outcome unit $i$ and each of the $J$ intervention units. $\mathbf{A}$ denotes the treatment status (or intervention status) vector. In particular, $\mathbf{A}=(A_{1}\dots A_{j}\dots A_{J})\in\{0,1\}^{J}$, where 1 corresponds to treating a particular power plant and 0 corresponds to not treating. Outcomes will depend on the exposure mapping (Aronow & Samii, 2017), which is given by $\bar{A}_{i}=\frac{1}{J}\sum_{j=1}^{J}H_{ij}A_{j}$— that is, the linear combination of treatments at the intervention unit level weighted by the strength of the connection between the outcome unit $i$ and intervention unit $j$ (interference map). Let $Y_{i}(\bar{a}_{i})$ denote the potential outcome for unit $i$ under exposure level $\bar{a}_{i}$ and $Y_{i}$ denote the observed outcome for outcome unit $i$. Note that, in this work, we will consider smaller values of $Y_{i}$ as desirable—e.g., a smaller number of IHD hospitalizations. Let the policy function for intervention unit $j$ be given by $\pi_{j}:\eta(\mathbf{X}^{out}_{i},\mathbf{H}_{j})\mapsto\Pi_{j}\subseteq[0,1]$, where $\eta$ is a functional of $\mathbf{X}^{out}_{i},\mathbf{H}_{j}$. Since we are interested in the health effects of some policy $\pi=(\pi_{1},\dots,\pi_{J})\in\Pi=[0,1]^{J}$, let $Y_{i}(\pi)$ be the potential outcome under policy $\pi$ for outcome unit $i$, similar to how ‘$Y(\rho)$’ is defined in Section 2 of Qiu et al. (2022). Finally, let $e_{j}$ be the propensity score for intervention unit $j$, or
$e_{j}=\mathbb{P}(A_{j}=1\mid\mathbf{X}^{int}_{j}).$
Objective.
@@
The objective is to determine the optimal treatment policy $\pi^{*}\in\Pi=[0,1]^{J}$ which minimizes $\mathbb{E}[Y_{i}(\pi)]$, the mean outcome under policy $\pi$— e.g. the treatment policy that yields the smallest number of IHD hospitalizations.
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi}\mathbb{E}[Y_{i}(\pi)]$ (1)
@@ 2.2 Nonparametric Identification
We state the identification assumptions required to identify our estimand Equation 1 from data.
@@
  1. (A1)
    Consistency of outcome units: $Y_{i}=Y_{i}(\bar{a}_{i})=Y_{i}(\mathbf{A})$.
  2. (A2)
    Positivity: $P(\bar{a}_{i}\mid\mathbf{X}^{out}_{i})>0$.
  3. (A3)
    Unconfoundedness and Intervention Unit Unconfoundedness: $Y_{i}(\bar{a}_{i})\perp\bar{a}_{i}\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J}$.
  4. (A4)
    Intervention covariates are independent of potential outcomes given the exposure mapping $\bar{a}_{i}$ and outcome covariates: $Y_{i}(\bar{a}_{i})\perp\mathbf{\mathbb{X}}_{1:J}\mid\mathbf{X}^{out}_{i},\bar{a}_{i}$.
Assumptions (A1)-(A3) are adaptations of the standard consistency, positivity, and ‘no unmeasured confounding’ assumptions in causal inference to our BNI setting. Note that the second equality of (A1) is made since $\mathbf{H}$ cannot be controlled and we only have control over $\mathbf{A}$. Assumption (A4) encodes domain knowledge from air pollution policy: outcome unit potential outcomes are determined by aggregate exposure levels $\bar{A}_{i}$ rather than individual treatment values of intervention units $A_{j}$. Thus, when learning the distribution of $Y_{i}(\bar{a}_{i})$ given $\bar{A}_{i}$ and $\mathbf{X}^{out}_{i}$, it is not necessary to condition on $\mathbf{\mathbb{X}}_{1:J}$. While this assumption is not strictly necessary, it enables simplifications of the proposed modeling approaches by leveraging domain knowledge.
Proposition 2.1 (Identification).
Under (A1)-(A4), we have
$\mathbb{E}[Y_{i}(\bar{a}_{i})]=\mathbb{E}\big{[}\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]\big{]}$ (2)
@@
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi}\mathbb{E}_{n}\big{[}\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]\big{]}$ (3)
Note that, in the real world, we may have cost constraints—e.g., a total budget for installation cost of scrubbers. Let $C$ denote the total cost constraint and for $j\in[J],$ let $c_{j}=\mbox{Cost}(A_{j}=1)-\mbox{Cost}(A_{j}=0)$ denote the cost difference between treating and non treating the intervention unit $j$. To enable identification of cost-constrained optimal policies, we require these costs $c_{j}$ to be known.
Cost Assumption
  1. (A5)
@@
Corollary 1 (Cost-constrained Optimal Policy).
Assume (A1)-(A4) and (A5). Then, the optimal cost constrained optimization problem is given by:
[…]

### v3

@@ Abstract
The substantial effect of air pollution on cardiovascular disease and mortality burdens is well-established. Emissions-reducing interventions on coal-fired power plants—a major source of hazardous air pollution—have proven to be an effective, but costly, strategy for reducing pollution-related health burdens. Targeting the power plants that achieve maximum health benefits while satisfying realistic cost constraints is challenging. The primary difficulty lies in quantifying the health benefits of intervening at particular plants. This is further complicated because interventions are applied on power plants, while health impacts occur in potentially distant communities, a setting known as bipartite network interference (BNI). In this paper, we introduce novel policy learning methods based on Q- and A-Learning to determine the optimal policy under arbitrary BNI. We derive asymptotic properties and demonstrate finite sample efficacy in simulations. We apply our novel methods to a comprehensive dataset of Medicare claims, power plant data, and pollution transport networks. Our goal is to determine the optimal strategy for installing power plant scrubbers to minimize ischemic heart disease (IHD) hospitalizations under various cost constraints. We find that annual IHD hospitalization rates could be reduced in a range from 23.37–55.30 per 10,000 person-years through optimal policies under different cost constraints.
keywords:
Environmental Policy; Ischemic Heart Disease; Power Plants; Optimal Treatment Regime
@@ 1 Introduction
From a statistical perspective, policy learning methods for emissions-reducing policies must accommodate two key problem features that differ significantly from more conventional causal inference settings. First, intervention strategies are implemented in coal-fired power plants (“intervention units”), but health impacts are measured in surrounding communities (“outcome units”). This creates a so-called bipartite data structure. Second, the complex process in which airborne pollutants react in the atmosphere and are transported (e.g., by the wind) means that intervening at a single power plant can potentially affect health in many distant communities, resulting in a network of connections between power plants and proximal communities. In causal inference, the presence of such connections creates interference, where one unit’s treatment can spill to (possibly many) untreated units and can affect their potential outcomes. The aggregate of these two features has been described as bipartite network interference (BNI) (Zigler and Papadogeorgou, 2021). Although a small number of papers have proposed approaches for conducting causal inference under BNI, optimal policy learning methods under BNI have not been explored to date.
@@ 1.1 Related Work
The literature on causal analysis under BNI is relatively sparse. Here, we highlight primary works in BNI, non-BNI, and spatial settings. The causal BNI literature began with Zigler and Papadogeorgou (2021), introducing estimators for causal effects under partial interference, adapting standard inverse probability weighted estimators to this simplified setting. The BNI literature expanded to accommodate more general causal estimands under experimental (Pouget-Abadie et al., 2019; Doudchenko et al., 2020) and observational settings (Zigler and Papadogeorgou, 2021; Zigler et al., 2023; Chen et al., 2024b), including average and heterogeneous treatment effects. More recent work considers the estimation of causal effects from longitudinal data (Chen et al., 2024a; Song and Papadogeorgou, 2024). To the best of our knowledge, no work has considered the problem of policy learning in the context of BNI. The closest related works consider policy learning under interference in non-BNI settings (Su et al., 2019; Viviano and Rudder, 2020; Viviano, 2024; Zhang and Imai, 2024).
From a methodological perspective, our proposed approach significantly differs from recent works in the field of policy learning and policy optimization. Viviano and Rudder (2020) propose experimental designs to study policy learning under interference when the interference network is unknown, but data are organized into a finite number of large clusters. Viviano (2024) proposes policy learning methods under an experimental setup for networks with an anonymous, limited degree of interference. Zhang and Imai (2024) propose policy learning methods under clustered interference, in which the interference is restricted to a finite cluster size $m$. To overcome high variance during estimation, they propose a semiparametric heterogeneous additive model, similar to ours, that practically restricts the level of interference. Furthermore, Su et al. (2019) proposes estimation and inference for network models under anonymous interference using a similar additive heterogeneous model to our approach without clustering assumptions. However, Su et al. (2019)’s approach assumes binary connections between units and a model setup that is specific to non-BNI settings. Methodologically, these works are significantly different from our proposed approach and our scientific application since (i) the BNI setting involves separate intervention units and outcome units meaning the estimands, treatment policy, and models are different as well, (ii) finite dependence clustered network interference is not practical for our setting as interference may extend past $m$ finite units, even if the spillover might be weak, and (iii) the exposures experienced by outcome units are variable based on climate patterns, and we wish to incorporate this information into our model.
From a theoretical perspective, we employ the multiplier central limit theorem to derive inferential results as Su et al. (2019) and spatial mixing along the lines of Jenish and Prucha (2009) to derive regret bounds for treatment effect functions under arbitrary interference. Jenish and Prucha (2009) derive general asymptotic results for random fields but do not consider causal estimation. We assume a subset of the infinite-sampling regime and mixing from Jenish and Prucha (2009), and employ an independent blocking technique from Bernstein (1927) adapted to spatial domains to derive regret bounds. Spatial literature tends to assume finite $m$-dependence or more general dependence via spatial mixing. Pinkse et al. (2007) utilizes independent blocking for spatial data but does not present the assumption set in an interpretable or verifiable manner and does not apply their methods to causal estimation. Jenish and Prucha (2012) consider asymptotics for near-epoch dependent fields, a popular notion in time-series literature, and prove results on $MA(\infty)$ models, but do not consider causal estimation. Guan et al. (2021) consider learning treatment regimes under spatial data for malaria control, but consider limited $m$-dependence for interference and non-bipartite settings. Other works have considered policy learning problems under safety or budget constraints, but operate under non-BNI and non-interference settings (Laber et al., 2018; Qiu et al., 2022; Xu et al., 2024). Adapting such work to BNI settings is left for future work.
@@ 1.2 Contributions and Organization
In this paper, we develop and apply methods to learn optimal power plant scrubber installation policies to minimize hospitalization burdens for ischemic heart disease (IHD) under cost constraints. Our approach will leverage a rich dataset of Medicare claims, power plant locations and features, power plant pollution transport networks, community features, and scrubber installation costs for the eastern U.S. Building methodologically on the work of Robins (2004), we propose Q-Learning and A-Learning methods for policy learning under arbitrary BNI. In contrast to much of the existing work on interference, our approach enables estimation and inference for the average treatment effects of intervening on intervention units without the need for (1) clustering assumptions or (2) causal estimands that might suffer of subjective choices of the exposure mapping such as spillover or direct effects (Hudgens and Halloran, 2008; Bargagli-Stoffi et al., 2025). We derive the asymptotic properties of these methods and demonstrate their finite sample performance via simulation studies, finding robustness of A-Learning to model misspecification in terms of estimation and inference, compared to Q-Learning. We use the proposed methods to determine the power plant scrubber installation strategy that would optimally reduce IHD hospitalization outcomes within the United States, under cost constraints, finding that annual IHD hospitalizations could be reduced in a range from 15,025 (23.37 per 10,000 person-years) to 38,372 (55.30 per 10,000 person-years) based on different financial budget constraints.
The rest of the paper is organized as follows: Section 2 introduces our Q- and A-Learning methods and their theoretical properties. Section 3 details an extensive semi-simulated Monte Carlo study carried out to assess performance of the methods in finite samples. In Section 4, we present the details and results of our scrubber installation policy analysis. Finally, we end with discussion and limitations in Section 5. Proofs are deferred to Appendix Appendix J.
@@ 2.1.1 Setup and Notation
Suppose we have $J$ intervention units indexed by $j\in[J]$ and $n$ outcome units indexed by $i\in[n]$. Let $\mathbf{X}^{out}_{i}$ denote the vector of covariates corresponding to outcome unit $i$, and $\mathbf{X}^{int}_{j}$ denote the vector of covariates corresponding to intervention unit $j$. $\mathbf{\mathbb{X}}_{1:J}$ will denote the covariates for all intervention units or $\mathbf{\mathbb{X}}_{1:J}=\{\mathbf{X}^{int}_{j}\}_{j\in[J]}$. $\mathbf{H}\in\mathbb{R}^{n\times J}$ or $\{\mathbf{H}_{ij}\}$ denotes the ‘interference map’ or bipartite adjacency matrix. The elements of $\mathbf{H}$ correspond to the strength of pollution transport from a particular intervention unit $j$ to each outcome unit $i$. $\mathbf{H}_{j}^{\top}=(\mathbf{H}_{1j}\dots\mathbf{H}_{ij}\dots\mathbf{H}_{nj})^{\top}$ denotes a particular column of the interference map. Similarly, $\mathbf{H}_{i}=(\mathbf{H}_{i1}\dots\mathbf{H}_{ij}\dots\mathbf{H}_{iJ})$ denotes a particular row of the interference map—that is, $\mathbf{H}_{i}$ represents the strength of connections between a particular outcome unit $i$ and each of the $J$ intervention units. In our scientific application, $\mathbf{H}$ is derived from meteorological variables (primarily wind patterns), measured with high accuracy from the National Oceanic and Atmospheric Administration (NOAA) Air Resources Laboratory, aggregated over space and time (Henneman et al., 2019). $\mathbf{A}$ denotes the treatment status (or intervention status) vector. In particular, $\mathbf{A}=(A_{1}\dots A_{j}\dots A_{J})\in\{0,1\}^{J}$, where 1 corresponds to treating a particular power plant and 0 corresponds to not treating. Outcomes will depend on the exposure mapping (Aronow and Samii, 2017), which is given by $\bar{A}_{i}=\frac{1}{J}\sum_{j=1}^{J}H_{ij}A_{j}$— that is, the linear combination of treatments at the intervention unit level weighted by the strength of the connection between the outcome unit $i$ and intervention unit $j$ (interference map). Let $Y_{i}(\bar{a}_{i})$ denote the potential outcome for unit $i$ under exposure level $\bar{a}_{i}$ and $Y_{i}$ denote the observed outcome for outcome unit $i$. Note that, in this work, we will consider smaller values of $Y_{i}$ as desirable—e.g., a smaller number of IHD hospitalizations. Let the policy function be given by $\pi:\eta(\mathbf{X}^{out}_{i},\mathbf{H}_{j})\mapsto[0,1]$, where $\eta$ is a functional of $\mathbf{X}^{out}_{i},\mathbf{H}_{j}$. Additionally define $e_{j}$ as the propensity score for intervention unit $j$, or $e_{j}=\mathbb{P}(A_{j}=1\mid\mathbf{X}^{int}_{j})$. Notationally, $\mathbb{E}$ will denote the mean with respect to any variables not explicitly conditioned on including $\mathbf{X}^{out}_{i},\mathbf{X}^{int}_{j},\mathbf{H}_{i}$ and $Y(\bar{a}_{i})$, and $\mathbb{E}_{\pi}$ will mean the expectation taken under policy $\pi$. $\mathbb{E}_{n}$ denotes the empirical mean as in the empirical process literature Kosorok (2008). $\lesssim$ will mean $\leq$ up to a finite constant.
2.1.2 Objective
@@
The objective is to determine the optimal treatment policy $\pi^{*}\in\Pi$ for some policy class $\Pi$, which minimizes $\mathbb{E}_{\pi}[Y_{i}(\bar{a}_{i})]$, the mean outcome under policy $\pi$— e.g. the treatment policy that yields the smallest number of IHD hospitalizations. The causal estimands depend on a realization of $\mathbf{H}$ and $n,J$. As done in Ogburn et al. (2024) and preceding works, the causal estimands and estimators condition on $\mathbf{H}$ and $n,J$.
$\pi^{*}=\arg\min_{\pi\in\Pi}\mathbb{E}_{\pi}[Y_{i}(\bar{a}_{i})]$ (1)
@@ 2.2 Nonparametric Identification
We state the identification assumptions required to identify our estimand Eq. 1 from data:
@@
  1. (I1)
    Consistency of outcome units: $Y_{i}=Y_{i}(\bar{a}_{i})$.
  2. (I2)
    Positivity: $P(\bar{a}_{i}\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J},\mathbf{H}_{i})>0$.
  3. (I3)
    Unconfoundedness: $Y_{i}(\bar{a}_{i})\perp\bar{a}_{i}\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J},\mathbf{H}_{i}$.
  4. (I4)
    Intervention covariates are independent of potential outcomes given the exposure mapping $\bar{a}_{i}$ and outcome covariates: $Y_{i}(\bar{a}_{i})\perp\mathbf{\mathbb{X}}_{1:J}\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{a}_{i}$.
Assumptions (I1) – (I3) are adaptations of the standard consistency, positivity, and ‘no unmeasured confounding’ assumptions in causal inference to our BNI setting. (I4) encodes domain knowledge from air pollution policy: outcome unit potential outcomes are determined by aggregate exposure levels $\bar{A}_{i}$ rather than individual treatment values of intervention units $A_{j}$.
Remark 1.
Assumption (I1) is made in terms of our exposure level $\bar{a}_{i}$, not at the intervention level $A_{j}$. Under this formulation, it is possible for two treatments $\mathbf{A}^{(1)}=(A_{1}^{(1)}\dots A_{J}^{(1)})\neq\mathbf{A}^{(2)}=(A_{1}^{(2)}\dots A_{J}^{(2)})$ to yield the same exposure level $\bar{a}_{i}$. This does not threaten the validity of our consistency assumption and results developed below, since we reason at the exposure level.
Remark 2.
We comment further on (I4): Reiterating our justification above, such an assumption is reasonable in our setting since health outcomes are impacted by total pollution levels, not individual power plant treatment statuses. Further, the predominant mechanism through which power plant covariates affect our health outcomes is air pollution levels. To account for additional confounders such as socioeconomic factors, we additionally adjust for $\mathbf{X}^{out}_{i}$ in the model. Thus, when learning the distribution of $Y_{i}(\bar{a}_{i})$ given $\bar{A}_{i}$ and $\mathbf{X}^{out}_{i}$, it is not necessary to condition on $\mathbf{\mathbb{X}}_{1:J}$.
Proposition 1 (Identification).
Under (I1) – (I4), we have
$\mathbb{E}[Y_{i}(\bar{a}_{i})]=\mathbb{E}\big{[}\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i}=\bar{a}_{i}]\big{]}$ (2)
@@
$\pi^{*}=\arg\min_{\pi\in\Pi}\mathbb{E}_{n}\big{[}\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i}=\bar{a}_{i}]\big{]}$ (3)
Note that, in the real world, we may have cost constraints—e.g., a total budget for installation cost of scrubbers. Let $C$ denote the total cost constraint and for $j\in[J],$ let $c_{j}=\mbox{Cost}(A_{j}=1)-\mbox{Cost}(A_{j}=0)$ denote the cost difference between treating and non treating the intervention unit $j$. To enable identification of cost-constrained optimal policies, we require these costs $c_{j}$ to be known: $c_{j}\geq 0$.
  1. (C1)
@@
Proposition 2 (Cost-constrained Optimal Policy).
Assume (I1) – (I4) and (C1). Then, the optimal cost constrained optimization problem is given by:
$\pi^{*}=\arg\min_{\pi\in\Pi}\quad$ $\mathbb{E}_{n}[\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i}=\bar{a}_{i}]]\quad\mbox{ s.t. }\sum_{j=1}^{J}c_{j}\pi_{j}\leq C\mbox{ }$ (4)